OpenStudy (anonymous):

1. What is the name of the first number in an ordered pair? A. w coordinate B. x coordinate C. y coordinate D. z coordinate 2. What is the name of the second number in an ordered pair? A. w coordinate B. x coordinate C. y coordinate D. z coordinate 3. What is STEP 1 of graphing a line using x and y tables? A. Choose 5 x values B. Choose 5 y values C. Graph points

OpenStudy (anonymous):

1.B 2.Y

OpenStudy (anonymous):

2.c

OpenStudy (anonymous):

B C and C

OpenStudy (anonymous):

4. What is STEP 2 of graphing a line with x and y tables? A. Choose 5 x values B. Create the ordered pairs by plugging the x-values you chose in step 1 into the given equation, and evaluating. C. Graph it

OpenStudy (anonymous):

1. B 2. C 3. C

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Step 2 is B

OpenStudy (anonymous):

7. Given 3x - y = 15, find y when x = -4 A. 27 B. 3 C. -3 D. -27 8. Given 3x - y = 15, find y when x = -2 A. -21 B. -9 C. 9 D. 21 9. Given 3x - y = 15, find y when x = 2 A. -21 B. -9 C. 9 D. 21 10. Given 3x - y = 15, find y when x = 4 A. -27 B. -3 C. 3 D. 27

OpenStudy (anonymous):

math sucks i cant never get it

OpenStudy (anonymous):

a a b b

OpenStudy (anonymous):

4. B 7. -27 8. -21 9. -9 10. -3

OpenStudy (anonymous):

thank you so much i wrote it all down for reference.

OpenStudy (anonymous):

yup(:

OpenStudy (anonymous):

thanks both of you!